a carnot cycle operate as a heat engine between two bodies of equal heat capacities until their temperature becomes equal if initial temperature of bodies s T1 ,T2 then t their common temperature
1
Expert's answer
2019-04-02T10:56:24-0400
"1-\frac{\delta Q_C}{\delta Q_H}=1-\frac{T_C}{T_H} (1)"

where


"\delta Q_C=C_pdT_c"

"\delta Q_H=C_pdT_H"

Integrating that condition through the whole process


"T=\sqrt{T_1T_2}"

Answer:


"T=\sqrt{T_1T_2}"